OpenStudy (anonymous):

Q 7) Find the missing term in the following table: x 1 2 3 4 5 6 7 f(x) 2 4 8 - 32 64 128

OpenStudy (anonymous):

16

OpenStudy (anonymous):

it can be clearly seen that f(x)=2^x hence f(4)=2^4

OpenStudy (anonymous):

solve it by interpolation

OpenStudy (anonymous):

i know the backward and forward interpolation formulas but what is the interpolation of central number